Write down Cp/Cv ratio for triatomic gas


1
Expert's answer
2021-01-26T08:16:57-0500

"\frac{C_p}{C_V}=\gamma,"

"\gamma=\frac{i+2}{i},"

"i=6,"

"\gamma=\frac{6+2}{6}=\frac 86=\frac 43,"

"\frac{C_p}{C_V}=\frac 43\approx 1.33."